OpenStudy (anonymous):

Solve for y. -1 + y 8 < 15

OpenStudy (solomonzelman):

You didn't post the equation.

OpenStudy (anonymous):

<

OpenStudy (anonymous):

is it -1+y8<15?

OpenStudy (solomonzelman):

\[-1+y+8<15\] right?

OpenStudy (anonymous):

yes it is -1+y8<15 @catjean

OpenStudy (solomonzelman):

what would y8 mean? y*8 y^8 or what?

OpenStudy (anonymous):

8y?

OpenStudy (solomonzelman):

\[-1+8y<15\]

OpenStudy (anonymous):

y over 8

OpenStudy (solomonzelman):

\[-1+\frac{y}{8}<15\]

OpenStudy (anonymous):

yes @solomonzelman

OpenStudy (solomonzelman):

\[\color{red} {+1}-1+\frac{y}{8}<15\color{red} {+1}\]\[\frac{y}{8}<16\]\[\frac{y}{8}\color{red} {\times 8}<16\color{red} {\times 8}\]\[\color{blue} {y~=~?}\]

OpenStudy (anonymous):

what does it =?

OpenStudy (anonymous):

Are we looking for the solution set?

OpenStudy (solomonzelman):

Question mark, I am asking you to tell me what the y is, based on the rest of the work.

OpenStudy (anonymous):

were solving for y

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

See what I was doing in each step?

OpenStudy (anonymous):

so y<128?